# Scenario analysis with multivariate Bayesian machine learning models

### v1

@@ 4. Empirical Applications
We employ the proposed framework in three related yet distinct applications. First, we use the annual stress test scenarios conducted by the Federal Reserve System and compute conditional forecasts using soft constraints on several variables. Second, we study the implications of varying financial conditions (imposing hard constraints over a period) on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock and gauge the role of spillovers and spillbacks to several other major economies. Across these applications we mainly compare a homoskedastic (BART-hom, setting $s_{t}=1$ for all $t$) and heteroskedastic BART implementation (BART-het) with a heteroskedastic BVAR (BVAR-het), both of which feature the outlier specification. Due to our sample featuring the Covid-19 pandemic observations we disregard the homoskedastic BVAR version (see Lenza and Primiceri, 2022, for a discussion).
@@ 4.1. Stress testing scenarios for the US economy
Major differences across model implementations emerge when considering measures of economic activity and payroll employment. While all three models agree about the fact that the adverse scenario leads to a significant economic contraction, the heteroskedastic BART model suggests somewhat less severe magnitudes of this downturn. This can be explained by noting that for this specification, the algorithm decides to classify several observations (that are otherwise informative about directional movements of variables when assuming homoskedasticity) as noisy or even outliers. We note that BART, due to the way how tree-based approaches fit data, is capable of dealing with outliers and heteroskedastic data features in the conditional mean function by design (see Huber et al., 2023; Clark et al., 2023, for discussions), even when assuming homoskedastic reduced form errors. But in this case it classifies several observations instead as noise rather than signal.
@@ 4.2. Financial conditions in the US and tail risk scenarios
In this empirical application, we restrict our sample from 1976Q1 to 2017Q4 and consider the period from 2018Q1 until 2019Q1 as a laboratory to assess nonlinearities in the relationship between economic variables and financial conditions. For this application we use the homoskedastic BART model as the pandemic observations are excluded from our dataset. We investigate nonlinear patterns of macroeconomic risk, which, following the “growth-at-risk” approach of Adrian et al. (2019), is defined as the predictive quantiles of some variable of interest at a pre-defined probability level (in line with value-at-risk, VaR, in finance). We pick this period because the information set already contains the global financial crisis (and the model thus had the opportunity to learn from this severe financial episode), and because this “holdout sample” otherwise coincides with a comparatively eventless period. We impose hard constraints on the future path of the National Financial Conditions Index (NFCI) and trace the effects of these scenarios on several macroeconomic variables. The scenarios are defined to reflect an increase of the NFCI ranging from approximately $1$, $3$ and $6$ unconditional standard deviations (reflecting tighter financial conditions) in 2018Q1, which we implement by placing these values as hard restriction on the NFCI in that quarter. From 2018Q2 onward we leave the respective future path unrestricted.
We investigate growth-at-risk (quantiles of real GDP), inflation-at-risk (quantiles of PCE inflation) and labor-at-risk (quantiles of growth in payroll employment) as our objects of interest (see Adams et al., 2021; Botelho et al., 2024; Clark et al., 2024; Lopez-Salido and Loria, 2024, for related papers). The resulting conditional forecast distributions (density estimates) are shown in Figure 4. Computing the difference between these conditional scenario distributions and the unconditional one yields an estimate of the unorthogonalized IRF (UIRF). In the spirit of Equation (10), we have $\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1:h})-\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I})$, which is the counterfactual difference between the scenario defined by the restrictions in the $\mathcal{C}_{h}$’s and the unconditional forecast. Different to a simulation of an individual structural shock, the UIRF reflects a likely combination of structural shocks that increases the NFCI by a predefined amount, see also Crump et al. (2025).
@@
The resulting predictive distributions exhibit non-Gaussian features, chief among them being particularly heavy tails and skewness (see also Clark et al., 2023, for a related but simplified scenario analysis in this context). In addition, there are hints of multimodality as the assumed values for the NFCI in 2018Q1 turn more extreme, which relates to the discussions in Adrian et al. (2021). For instance, while most mass of the predictive distribution of payroll employment growth, for the extreme stress scenario, is centered on about $-3$ percent, there is a smaller local peak at values of $-5$ percent. A similar pattern is visible for PCE inflation.
@@ 4.3. Spillovers and spillbacks of US financial shocks
This application estimates the effect of a financial shock in the US and traces its effects through the domestic economy, but also captures spillovers and spillbacks to and from several other economies. We use a slightly different dataset in this case, which drops several of the domestic indicators, but adds bilateral exchange rates with the EA and Japan alongside real GDP for the EA, Japan and the UK.
Following the SVAR literature, we assume that the reduced form shocks are linked to the structural shocks as follows, $\bm{\epsilon}_{t}=\bm{A}_{0}^{-1}\bm{u}_{t}$. This implies that $\bm{\Sigma}=\bm{A}_{0}^{-1}\bm{A}_{0}^{-1}{{}^{\prime}}$, and we identify $\bm{A}_{0}^{-1}$ using a set of zero impact restrictions imposed with a Cholesky decomposition of the form $\bm{\Sigma}=\bm{D}\bm{D}^{\prime}$. That is, $\bm{A}_{0}^{-1}=\bm{D}$ and $\partial\bm{y}_{\tau}/\partial u_{j\tau}=\bm{D}\bm{e}_{j}^{\prime}=\bm{\delta}_{0}$ and we orthogonalize the structural shocks (different to the application in Section 4.2 which dealt with unorthogonalized shocks that resulted in an increase of the NFCI). We simulate different shock sizes of different signs with $d\in\{-3,-1,1,3,6\}$. Different to conventional linear SVAR frameworks, our approach allows to assess nonlinearities of responses with respect to different signs and magnitudes of a shock with GIRFs. Such asymmetries and related nonlinearities have recently gained attention both in a VAR and local projection context, see, e.g., Mumtaz and Piffer (2022); Carriero et al. (2023); Forni et al. (2024); Hauzenberger et al. (2024a).
To identify the financial shock, we place timing-restrictions on the contemporaneous impulse responses as described above. This is operationalized with a specific ordering of the quantities in the vector $\bm{y}_{t}$. Specifically, we structure this vector such that all slow moving domestic and foreign macroeconomic variables come first (which imposes zero restrictions on impact). These variables are then followed by the excess bond premium (EBP, Gilchrist and Zakrajšek, 2012), and all fast moving variables such as those capturing the financial economy. We interpret the orthogonalized innovation of the EBP equation as the financial shock, similar to Barnichon et al. (2022). In a multicountry context, Huber et al. (2024) use an identification scheme virtually identical to ours.
@@
We present selected results for key macroeconomic and financial variables in Figure 5, additional empirical results can be found in the Appendix. While our framework allows to compute GIRFs for each period in our sample, we focus on time averages in the results that follow. The rows in the figure show different subsets of the same set of results, structured such that the shocks of different signs and sizes can be compared with ease. That is, for visualization purposes, we typically rescale the GIRFs to a common range by computing “normalized” GIRFs, $\tilde{\bm{\lambda}}_{\tau,h}^{m,d}=\bm{\lambda}_{\tau,h}^{(m,d)}/d$. Note that for linear VARs, such scalings do not make a difference because the responses are symmetric (see the additional results for the linear BVAR in the Appendix). For nonlinear models, this is not necessarily the case and allows for a visual inspection of asymmetries.
@@ 5. Conclusions
This paper presents a robust and unified methodology for conducting scenario analysis in multivariate macroeconomic settings, accommodating nonlinearities and unknown functional forms of conditional mean relationships. The proposed methods are applicable not only in traditional nonlinear frameworks, such as variants of threshold or time-varying parameter models, but also to more recently developed models incorporating machine learning techniques. By extending conditional forecasts and GIRFs to a flexible, nonparametric framework, we address some limitations of traditional linear and parametric models in generating various types of counterfactual analyses. Specifically, we explore the application of conditional forecasts in a nonlinear setting by leveraging the properties of conditionally Gaussian errors. Furthermore, we demonstrate how to derive dynamic causal effects in the form of GIRFs in a nonlinear context and quantify the contribution of specific transmission channels in the propagation of structural shocks. This approach is broadly applicable and computationally efficient, making it suitable for large-scale macroeconomic datasets.
Empirical applications, focusing on the use of BART as an example of nonparametric modeling of the conditional mean function, underscore the critical role of nonlinearities and heteroskedasticity in shaping macroeconomic dynamics. For instance, scenario analysis based on Federal Reserve stress tests reveals differences between linear and nonlinear models in forecasting economic contractions and recoveries. Similarly, our growth-at-risk application demonstrates how nonlinearities influence the distribution of macroeconomic risks under financial stress, particularly in amplifying downside risks. Finally, the analysis of financial spillovers reveals significant asymmetries in the transmission of shocks and the influence of international linkages on domestic outcomes. Our results indicate that incorporating nonlinearities in conditional forecasts and scenario analysis provides a richer understanding of risk propagation and policy effects. By addressing limitations in traditional linear approaches and offering a flexible tool for analyzing complex economic relationships, this paper contributes to the literature on macroeconomic forecasting, risk assessment, and policy evaluation.

### v2

@@ 4. Empirical Applications
We employ the proposed framework in three related yet distinct applications. First, we use the annual stress test scenarios conducted by the Federal Reserve System and compute CFs using soft constraints on several variables. Second, we study the implications of varying financial conditions (imposing a hard constraint on a period) on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock and gauge the role of spillovers and spillbacks to several other major economies. Across these applications we mainly compare a homoskedastic (BART-hom, setting $s_{t}=1$ for all $t$) and heteroskedastic BART implementation (BART-het) with a heteroskedastic BVAR (BVAR-het), both of which feature the outlier specification. Due to our sample featuring the Covid-19 pandemic observations we disregard the homoskedastic BVAR version (see Lenza and Primiceri, 2022, for a discussion).
@@ 4.1. Stress Testing Scenarios for the US Economy
Major differences across model implementations emerge when considering measures of economic activity and payroll employment. While all three models agree about the fact that the adverse scenario leads to a significant economic contraction, the heteroskedastic BART model suggests somewhat less severe magnitudes of this downturn. This can be explained by noting that for this specification, the algorithm decides to classify several observations (that are otherwise informative about directional movements of variables when assuming homoskedasticity) as more noisy, or even as outliers. We note that BART, due to the way how tree-based approaches fit data, is capable of dealing with outliers and heteroskedastic data features in the conditional mean function by design (see Huber et al., 2023; Clark et al., 2023, for discussions), even when assuming homoskedastic reduced form errors. But in this case it classifies several observations instead as noise rather than signal.
@@ 4.2. Financial Conditions in the US and Tail Risk Scenarios
In this empirical application, we restrict our sample from 1976Q1 to 2017Q4 and consider the period from 2018Q1 until 2019Q1 as a laboratory to assess nonlinearities in the relationship between economic variables and financial conditions. For this application we use the homoskedastic BART model as the pandemic observations are excluded from our dataset. We investigate nonlinear patterns of macroeconomic risk, which, following the “growth-at-risk” approach of Adrian et al. (2019), is defined as the predictive quantiles of some variable of interest at a pre-defined probability level (in line with value-at-risk, VaR, in finance). We pick this period because the information set already contains the global financial crisis (and the model thus had the opportunity to learn from this severe financial episode), and because this “holdout sample” otherwise coincides with a comparatively eventless period. We impose hard constraints on the future path of the National Financial Conditions Index (NFCI) and trace the effects of these scenarios on several macroeconomic variables.
The scenarios are defined to reflect an increase of the NFCI ranging from approximately $1$, $3$ and $6$ unconditional standard deviations (reflecting tighter financial conditions) in 2018Q1, i.e., in $\mathcal{C}_{1}$, which we implement by placing these values as hard restriction on the NFCI in that quarter. From 2018Q2 onward we leave the respective future path unrestricted, i.e., $\mathcal{C}_{h}=\varnothing$ for $h>1$. It is worth mentioning that in this specific case the two possible variants of CF as discussed in Section 2.2 are identical.
We investigate growth-at-risk (quantiles of real GDP), inflation-at-risk (quantiles of PCE inflation) and labor-at-risk (quantiles of growth in payroll employment) as our objects of interest (see Adams et al., 2021; Pfarrhofer, 2022; Botelho et al., 2024; Clark et al., 2024; Lopez-Salido and Loria, 2024, for related papers). The resulting CF distributions (density estimates) are shown in Figure 4. Computing the difference between these conditional scenario distributions and the unconditional one yields an estimate of the unorthogonalized IRF (UIRF). Reflecting Equation (10), we have $\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1})-\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I})$, which is the counterfactual difference between the scenario defined by the restriction in $\mathcal{C}_{1}$ and the unconditional forecast. Different to a simulation of an individual structural shock, the UIRF reflects a likely combination of structural shocks that increases the NFCI by the predefined amount, see also Crump et al. (2025).
@@
The resulting predictive distributions exhibit non-Gaussian features, chief among them being heavy tails and skewness (see also Clark et al., 2023, for a related but simplified scenario analysis in this context). In addition, there are hints of multimodality as the assumed values for the NFCI in 2018Q1 turn more extreme, which relates to the discussions in Adrian et al. (2021). For instance, while most mass of the predictive distribution of payroll employment growth, for the extreme stress scenario, is centered on about $-3$ percent, there is a smaller local peak at values of $-5$ percent. A similar pattern is visible for PCE inflation. These features can arise — even in one-step-ahead predictions and for a single restricted period — due, for example, to the initial conditions of the nonlinear unconditional mean function at the time the forecast is generated.
@@ 4.3. Spillovers and Spillbacks of US Financial Shocks
This application estimates the effect of a financial shock in the US and traces its effects through the domestic economy, but also captures spillovers and spillbacks to and from several other economies. We use a slightly different dataset in this case, which drops several of the domestic indicators, but adds bilateral exchange rates alongside real GDP for the EA, Japan and the UK. Following the linear SVAR literature, we assume that the reduced form shocks are linked to the structural shocks as $\bm{\epsilon}_{t}=\bm{B}_{0}^{-1}\bm{u}_{t}$. This implies that $\bm{\Sigma}=\bm{B}_{0}^{-1}\bm{B}_{0}^{-1}{{}^{\prime}}$ and that shocks impact the dynamic model linearly. Suppose the structural shock of interest is in the $j$th position of $\bm{u}_{t}$ and $\bm{u}_{-jt}=\bm{0}_{n-1}$, then the impact response is $d\bm{B}_{0}^{-1}\bm{e}_{j}^{\prime}=d\bm{\beta}_{0}$. That is, $\bm{\beta}_{0}$ is the $j$th column of the matrix $\bm{B}_{0}^{-1}$.
We identify $\bm{B}_{0}^{-1}$ using a set of zero impact restrictions imposed with a Cholesky decomposition of the form $\bm{\Sigma}=\bm{D}\bm{D}^{\prime}$ where $\bm{D}$ is lower triangular. That is, $\bm{B}_{0}^{-1}=\bm{D}$, and we orthogonalize the structural shocks (different to the application in Section 4.2 which dealt with unorthogonalized shocks that resulted in an increase of the NFCI). We simulate different shock sizes of different signs with $d\in\{-3,-1,1,3,6\}$. Different to conventional linear frameworks, our approach allows to assess nonlinearities of higher-order responses with respect to different signs and magnitudes of a proportional shock impact with GIRFs. Such asymmetries and related nonlinearities have recently gained attention both in a VAR and local projection context, see, e.g., Mumtaz and Piffer (2022); Carriero et al. (2023); Forni et al. (2024); Hauzenberger et al. (2024a).
To identify the financial shock, we place timing-restrictions on the contemporaneous impulse responses as described above. This is operationalized with a specific ordering of the quantities in the vector $\bm{y}_{t}$ — we structure this vector such that all slow moving domestic and foreign macroeconomic variables come first (which imposes zero restrictions on impact). These variables are then followed by the excess bond premium (EBP, Gilchrist and Zakrajšek, 2012), and all fast moving variables such as those capturing the financial economy. We interpret the orthogonalized innovation of the EBP equation as the financial shock, similar to Barnichon et al. (2022). In a multicountry context, Huber et al. (2024) use an identification scheme virtually identical to ours.
@@
We present selected results for key macroeconomic and financial variables in Figure 5, additional empirical results can be found in the Appendix. While our framework allows to compute GIRFs for each period in our sample, we focus on time averages in the results that follow. The rows in the figure show different subsets of the same set of results, structured such that the shocks of different signs and sizes can be compared with ease. That is, for visualization purposes, we typically rescale the GIRFs to a common range by computing “normalized” GIRFs, $\tilde{\bm{\delta}}_{\tau,h}^{(m,d)}=\bm{\delta}_{\tau,h}^{(m,d)}/d$. Note that for linear VARs, such scaling yields identical IRFs for all shock sizes (see the additional results for the linear BVAR in the Appendix). For nonlinear models, this is not necessarily the case and allows for a visual inspection of asymmetries.
@@ 5. Conclusions
This paper presents a robust and unified methodology for conducting scenario analysis in multivariate macroeconomic settings, accommodating nonlinearities and unknown functional forms of conditional mean relationships. The proposed methods are applicable not only in traditional nonlinear frameworks, such as variants of threshold or time-varying parameter models, but also to more recently developed models incorporating machine learning techniques. By extending CFs and GIRFs to a flexible, nonparametric framework, we address some limitations of traditional linear and parametric models in generating various types of counterfactual analyses. Specifically, we explore the application of CFs in a nonlinear setting by leveraging the properties of conditionally Gaussian errors. Furthermore, we demonstrate how to derive dynamic causal effects in the form of GIRFs in a nonlinear context and quantify the contribution of specific transmission channels in the propagation of structural shocks. This approach is broadly applicable and computationally efficient, making it suitable for large-scale macroeconomic datasets.
Empirical applications, focusing on the use of BART as an example of nonparametric modeling of the conditional mean function, underscore the critical role of nonlinearities and heteroskedasticity in shaping macroeconomic dynamics. For instance, a scenario analysis based on Federal Reserve stress test assumptions reveals differences between linear and nonlinear models in forecasting economic contractions and recoveries. Similarly, our growth-at-risk application demonstrates how nonlinearities influence the distribution of macroeconomic risks under financial stress, particularly in amplifying downside risks. Finally, the analysis of financial spillovers reveals significant asymmetries in the transmission of shocks and the influence of international linkages on domestic outcomes. Our results indicate that incorporating nonlinearities in CFs and scenario analysis provides a richer understanding of risk propagation and policy effects. By addressing limitations in traditional linear approaches and offering a flexible tool for analyzing complex economic relationships, this paper contributes to the literature on macroeconomic forecasting, risk assessment, and policy evaluation.